Project

General

Profile

NDOS Commissioning Meeting 2012/04/30

Attendance: Jon, Alex, Raphael, Susan, Nick, Denis, Athans, Minerba, Brian M., Ruth,

Agenda:

  1. Beam and Detector Update
  2. DataCheck Report
  3. Installation Plans for the week
  4. DAQ Update and Plans for the week
  5. DCS Update and Plans for the week
  6. AOB

Beam and Detector Report

  • All seems to going fine. 12 APDs are being actively purged and cooled, no problems seen so far (it's been about a week).
  • Mat did remote pedestal scans over the weekend using the VNC.

DataCheck Report:

  • None.

Installation plans for this week

  • The plan was to put new DCMs on
  • Leon will arrive tonight/tomorrow and Mat and Leon will begin installing
    • will start installing on downstream end of 3rd diblock and work their way upstream
    • 200 silicone-coated APDs
    • purge system is installed everywhere on NDOS
  • Mat:
    • end of next week: should all of diblock 3 instrumented and most of 2nd diblock too.
    • still waiting for new boxes to arrive, but we have enough FEBs.
    • should be able to do ~30/day

DAQ Update and Plans for next week

  • Peter: would like to do some more partitioning tests. Denis is currently working on making necessary modifications to DataLogger.
  • Kurt: would like to try new version of AppMgr

DCS Update and Plans for next week

  • Athans:
    • APD cooling monitoring SW (v2.0) is ready for deployment (new IOCs are running now).
    • Alarming discussion will be held on Thursday

AOB

  • Mat: some people have started to volunteer to do "skeleton shifts" (Brian M. has started). Athans: how will shifters be trained? Mat: will modify shift wiki with input from Brian.
  • Mat: we should do